A garden is 3 feet by 5 feet. How much will each side have to be increased in order to increase the area by 20 feet squared?
Answer by ankor@dixie-net.com(22740)

A garden is 3 feet by 5 feet. How much will each side have to be increased in order to increase the area by 20 feet squared?
:
Find original area: 3 * 5 = 15 sq/ft
Then new area = 15 + 20 = 35 sq/ft
:
Add x ft to each dimension to make it equal 35 sq/ft
(x+3)*(x+5) = 35
FOIL
x^2 + 5x + 3x + 15 = 35
:
x^2 + 8x + 15 - 35 = 0
:
x^2 + 8x - 20 = 0
Factors to
(x+10)(x-2) = 0
:
only the positive solution will make sense here
x = 2 ft added to each dimension
;
:
Check: (2+3) (2+5) = 35 sq/ft which is an increase of 20 sq/ft